The mean serum-creatinine level measured in 12 patients 24 hours after they received a newly proposed antibiotic was 1.2 mg/dL. 1. If the mean and standard deviation of serum creatinine in the general population are 1.0 and 0.4 mg/dL, respectively, then, using a significance level of 0.05, test whether the mean serum-cre level in this group is different from that of the general population. 2. What is the p-value for the test? 3. Suppose t1.52 and a on one-sample t test is performed based on seven subjects. What is the 4. Suppose the sample standard deviation of serum creatinine is now 0.6 mg/dL. Assume that th two-tailed p-value? standard deviation of serum creatinine is not known. Perform the p-value. hypothesis test and report a 5. Compute a two-sided 95% CI for the true mean serum-creatinine level. 6. How does your answer to part 5 relate to your answer to part 4?
Answer #1

1)

TS = (xbar - mu)/(sd/sqrt(n))

= (1.2 - 1)/(0.4/sqrt(12))

= 1.73205

critical value for 1-sample z-test is 1.96 for alpha = 0.05

since TS < critical value

we fail to reject the null hypothesis

2)

p-value= 2 P(Z > 1.73205) = 0.0833

3)

df = n-1 = 6

p-value = 2 P(t > 1.52) = 0.1793

Please post rest questions again